# zkPHIRE: A Programmable Accelerator for ZKPs over HIgh-degRee, Expressive Gates

### v1

@@ IV-B6 Interconnect and Memory
HyperPlonk’s stage-level data obliviousness enables static scheduling of computation and communication via a centralized controller. zkPHIRE comprises six major modules connected by a multi-channel shared bus, which is provisioned to meet peak data movement demands. During SumCheck and Wiring Identity, where bidirectional transfers occur between the SumCheck unit and Multifunction Forest, and one-way from PermCheck Generator to MSM unit, up to three channels are needed to avoid stalls. At our 294mm2 design point, the peak bandwidth requirement reaches 19 TB/s, comparable to state-of-the-art accelerators  [47, 28].
Each module includes local SRAM scratchpads to buffer intermediate data. Smaller buffers (6 MB) serve PermCheck Generator, MLE Combine, and the Multifunction Forest. Heavier modules—MSM (43 MB) and SumCheck (6 MB)—use highly banked SRAM to support multiple PEs and exploit data reuse. For instance, MSM stores elliptic curve points, while SumCheck stores frequently accessed MLE tiles. All modules connect to off-chip PHYs via the shared bus, which is coordinated to avoid contention and sustain peak throughput.
@@ V Methodology
[…]
The HyperPlonk CPU performance is measured on an AMD EPYC 7502 32-core processor running up to 3.35GHz, and the total die size is 296 mm2 [41, 5, 54]. We benchmark our SumCheck unit performance relative to prior work [12] and a 4-threaded CPU implementation. We then use workloads from prior work [9, 1] to model the performance of our full architecture vs. 32 CPU threads, up to problem sizes of $2^{30}$ nominal gates. We show these results in Section VI. We use a cycle-accurate simulator to model runtimes for MSMs. For SumCheck and tree-based computations, we model performance using analytical models, since these steps are data-oblivious. For sparse computations, we use the same workload statistics from prior work [13, 36, 66, 1].
@@ 5 $y$-gated Curve Check $(q_{\text{point}}\cdot y)\cdot(y^{2}-x^{3}-5)$
6 Incomplete Addition 1
$q_{\text{add-incomplete}}\cdot((x_{r}+x_{q}+x_{p})$
$\cdot(x_{p}-x_{q})^{2}-(y_{p}-y_{q})^{2})$
7 Incomplete Addition 2
$q_{\text{add-incomplete}}\cdot(y_{r}+y_{q})(x_{p}-x_{q})$
$(y_{p}-y_{q})(x_{q}-x_{r})$
8 Complete Addition 1
$q_{\text{add}}\cdot(x_{q}-x_{p})((x_{q}-x_{p})\lambda$
$(y_{q}-y_{p}))$
9 Complete Addition 2
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha)$
$(2y_{p}\lambda-3x_{p}^{2})$
10 Complete Addition 3
$q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$
$(\lambda^{2}-x_{p}-x_{q}-x_{r})$
11 Complete Addition 4
$q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$
$(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
12 Complete Addition 5
$q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$
$(\lambda^{2}-x_{p}-x_{q}-x_{r})$
13 Complete Addition 6
$q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$
$(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
@@ 17 Complete Addition 10 $q_{\text{add}}\cdot(1-x_{q}\cdot\gamma)(y_{r}-y_{p})$
18 Complete Addition 11
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$
$(y_{q}+y_{p})\delta)x_{r}$
19 Complete Addition 12
$q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$
$(y_{q}+y_{p})\delta)y_{r}$
HyperPlonk Polynomials [9]
20 Vanilla ZeroCheck
$(q_{L}w_{1}+q_{R}w_{2}-q_{O}w_{3}$
$q_{M}w_{1}w_{2}+q_{C})f_{r}$
21
Vanilla PermCheck
($\alpha$ is a scalar)
$(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}$
$N_{1}N_{2}N_{3}))f_{r}$
22 Jellyfish ZeroCheck
$(q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{3}+q_{4}w_{4}$
$q_{M_{1}}w_{1}w_{2}+q_{M_{2}}w_{3}w_{4}$
$q_{H_{1}}w_{1}^{5}+q_{H_{2}}w_{2}^{5}+q_{H_{3}}w_{3}^{5}$
$q_{H_{4}}w_{4}^{5}-q_{O}w_{5}$
$q_{ecc}w_{1}w_{2}w_{3}w_{4}+q_{C})f_{r}$
23
Jellyfish PermCheck
($\alpha$ is a scalar)
$(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}\cdot$
$D_{4}D_{5}-N_{1}N_{2}N_{3}N_{4}N_{5}))f_{r}$
20 OpenCheck
$y_{1}f_{r_{1}}+y_{2}f_{r_{2}}+y_{3}f_{r_{3}}$
$y_{4}f_{r_{4}}+y_{5}f_{r_{5}}+y_{5}f_{r_{5}}$
Figure 6: Speedups over 4-threaded CPU implementations of SumChecks on the polynomials in Table I. We pick a different optimal design across each bandwidth technology. Utilization across the design points shown
@@ VI-A1 Speedup over various gate types
Table I enumerates the polynomials we use in our evaluation. We first benchmark the performance of SumChecks on each polynomial by extending HyperPlonk’s existing code base [9]. We then run each polynomial on our programmable SumCheck unit. In this evaluation, we seek to optimize both performance and utilization across all polynomials. We can always allocate EEs and PLs to handle the worst-case polynomial degree; however, this would lead to underutilization on low-degree polynomials. In our design space search, we specify an area and bandwidth constraint and then use the following objective function:
$\min_{\text{design}}\;\lambda\cdot(1-f_{\text{util}}(u_{d,i}))+\lambda\cdot f_{\text{slowdown}}(s_{d,i})$
Here, $s_{d,i}$ refers to the slowdown of the $i^{th}$ polynomial for a given design $d$, relative to the fastest runtime for that polynomial in the area-constrained design space. Similarly, $u_{d,i}$ reflects the utilization for design $d$ on polynomial $i$. $f_{\text{slowdown}}$ and $f_{\text{util}}$ are aggregation functions; we employ geomean for slowdown and arithmetic mean for utilization. $\lambda$ controls the tradeoff between the minimizing slowdown and maximizing utilization. For this analysis, we focus on utilization, using $\lambda=0.8$. We compare the runtime against the CPU running each polynomial’s SumCheck on 4 threads, for which we estimate the total core area to be 37 mm2 in 7nm [41, 5, 54] and use that as our area constraint (after scaling from 22nm). Across various bandwidths, we find the optimal design point and report speedups and utilizations in Figure 6. For most designs, two EEs and five PLs is optimal.
We observe  50% utilization across benchmarks at each bandwidth, due to three main factors: (1) MLE Update units are inactive during the first round of SumCheck, which accounts for about half of total runtime; (2) lower-degree polynomials require fewer product lanes; and (3) repeating MLEs (e.g., $q_{add}$) reduce update modmul demand since they are reused across terms. Consequently, when optimizing for utilization, we observed a tendency to limit the number of EEs. Further parallelism and potentially higher utilization can be achieved by mapping multiple terms to EEs when possible, but this would introduce additional complexity in interconnect and increase bandwidth demand. These tradeoffs can be explored further in future work. In contrast, product lane modmuls are usually fully utilized, provided the polynomial degree and memory bandwidth are high enough. Despite moderate utilization, we achieve significant speedups over CPU baselines, reaching nearly 1000× at 1 TB/s of bandwidth.
VI-A2 High degree sweep
@@
We analyze the performance scaling of SumChecks as we increase polynomial degree. We sweep polynomial degrees ($d=2\ldots 30$) for $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+q_{c}$, as in HyperPlonk, and measure speedup over CPU. We use the same objective as before and pick a high-performance design under same area constraints (Figure 7). Our results show that low-degree polynomials require HBM-scale bandwidths to achieve $1000\times$ speedups, while higher-degree polynomials can reach similar speedups with DDR5-level bandwidths (e.g., 256 GB/s). Higher-degree polynomials in this scenario exhibit more computations on the same amount of data, which is why the disparity in speedups across bandwidths is less compared to low-degree polynomials.
Figure 8: Scheduler-induced behavior of SumCheck at fixed bandwidth and PL settings.
At fixed bandwidth, PE count, and EE-per-PE settings, increasing the number of PLs per PE improves performance roughly proportionally. Similar gains are observed when increasing EE count at fixed PLs. Interestingly, as shown in Figure 8, runtime increases in discrete jumps due to the graph decomposition (see Figure 2) the scheduler uses to map polynomials to EEs. For example, under 6 EEs, degree-1166 polynomials have 1 node in their schedule graph, while degree-771111 polynomials require 2. Each added node incurs a sharp runtime jump, as it introduces additional products across all MLEs and terms. Within each node cluster, runtime grows more gradually with degree due to early exit optimizations. These trends may help optimize the SumCheck unit when the degree distribution of polynomials is known, by minimizing high-degree outliers that cause runtime spikes.
@@ VI-A3 Comparison with Prior ASICs
Figure 9 details our speedups relative to a prior ASIC, zkSpeed. zkSpeed implements the SumChecks by using a custom unified core that maximizes immediate reuse. We also compare with zkSpeed+, which gives zkSpeed the benefits of pipelining MLE updates directly into the extensions and product computations (as we do in zkPHIRE). We assume the same modular multiplier area and 2 TB/s as zkSpeed. We use the same objective function from Section VI-A1 on the same “training set” of polynomials in Table I with a roughly iso-zkSpeed area constraint. We pick a 35.24 mm2 design to compare with zkSpeed’s 30.8 mm2 SumCheck and MLE Update area. Our design has 4.870 mm2 of SRAM, while zkSpeed used a much larger global SRAM for storing MLEs, hence we believe this comparison is fair.
@@
We find that, for ZeroCheck, because the Jellyfish polynomial is more complex, the $2\times$ is not sufficient to bring speedups over the vanilla counterpart. However, PermCheck’s increase in complexity is outweighed by a $2\times$ workload reduction. OpenCheck, which is the same between Jellyfish and Vanilla, exhibits $1.12\times,3.54\times$ and $6.22\times$ speedup for each gate reduction over zkSpeed+. Combining all three SumChecks, a Jellyfish gate count reduction by $4\times$ is sufficient to outperform vanilla polynomials on both zkSpeed+ and zkPHIRE. These gains are observed despite the inherent overheads of programmability in our accelerator.

### v2

@@ IV-B6 Interconnect and Memory
HyperPlonk’s stage-level data obliviousness enables static scheduling of computation and communication via a centralized controller. zkPHIRE comprises six major modules connected by a multi-channel shared bus, provisioned to meet peak data movement demands. During Wiring Identity, where bidirectional transfers occur between the SumCheck unit and Multifunction Forest, and one-way from Permutation Quotient Generator to MSM unit, up to three channels are needed to avoid stalls. At our 294mm2 design point, the peak bandwidth requirement reaches 19 TB/s, comparable to state-of-the-art accelerators [52, 28].
Each module includes local SRAM scratchpads to buffer intermediate data. Smaller buffers (6 MB) serve the Permutation Quotient Generator, MLE Combine, and the Multifunction Forest modules. Heftier modules—MSM (43 MB) and SumCheck (6 MB)—use highly banked SRAM to support multiple PEs and exploit data reuse. For instance, MSM stores elliptic curve points, while SumCheck stores frequently accessed MLE tiles. All modules connect to off-chip PHYs via the shared bus, coordinated to avoid contention and sustain peak throughput.
@@ V Methodology
We use Catapult HLS 2024 to generate the RTL for Montgomery multipliers with arbitrary primes (as in prior work [10, 9]) and fixed primes [39, 5], fully-pipelined PADD units, modular inverse units, and the SumCheck PEs. We use the same elliptic curve as HyperPlonk (BLS12-381), where all MLE datatypes (e.g., in the SumCheck unit) are 255-bit, and all elliptic curve points (e.g., in the PADD) are 381-bit. For logic synthesis, we use Design Compiler with TSMC 22nm. We explore the use of both arbitrary and fixed-prime modular multipliers (as done in [54]); the latter saves us roughly 50% on area and improves our computational density by roughly $2\times$. Our 255b modular multipliers are 0.478/0.264 mm2 (arbitrary/fixed); our 381b modular multipliers are 1.13/0.582 mm2 (arbitrary/fixed). We use Synopsys 22nm Memory Compiler for SRAM estimation. For SHA3, we use an IP block from OpenCores [47]. To match prior work, we use scale factors of $3.6$ for area and $3.3$ for power to scale to 7nm [10, 9, 8, 65, 66]. We use a 1GHz clock for zkPHIRE. For workload and sparsity, we use the same statistics from prior work [10, 38, 73, 31].
The HyperPlonk CPU performance is measured on an AMD EPYC 7502 32-core processor running up to 3.35GHz, and the total die size is 296 mm2 [44, 1, 58]. We benchmark the GPU performance on an NVIDIA A100 GPU with 40GB memory and 1.6 TB/s bandwidth [46]. We benchmark our SumCheck unit performance relative to prior work [9] and a 4-threaded CPU implementation. We then use workloads from prior work [6, 31] to model the performance of our full architecture vs. 32 CPU threads, up to problem sizes of $2^{30}$ nominal gates. We show these results in Section VI.
We use HLS tools to simulate core modules (modular multipliers, product lanes, extension engines, etc.) and use HLS-generated RTL to extract exact cycle counts, initiation intervals, and pipeline depths. We use cycle-accurate simulators with analytical models of bandwidth constraints to estimate runtimes for MSMs and for the traversal-dependent behavior of the Multifunction Forest [40]. For SumCheck, we integrate HLS-generated cycle counts into a higher-level model (the automated scheduler in Figure 2) and then simulate runtimes analytically (accounting for off-chip bandwidth). This approach is similar to prior work on accelerating cryptographic computing [9, 10, 52, 53].
@@ 5 $y$-gated Curve Check $(q_{\text{point}}\cdot y)\cdot(y^{2}-x^{3}-5)$
6 Incomplete Addition 1 $q_{\text{add-incomplete}}\cdot((x_{r}+x_{q}+x_{p})$ $\cdot(x_{p}-x_{q})^{2}-(y_{p}-y_{q})^{2})$
7 Incomplete Addition 2 $q_{\text{add-incomplete}}\cdot(y_{r}+y_{q})(x_{p}-x_{q})$ $(y_{p}-y_{q})(x_{q}-x_{r})$
8 Complete Addition 1 $q_{\text{add}}\cdot(x_{q}-x_{p})((x_{q}-x_{p})\lambda$ $(y_{q}-y_{p}))$
9 Complete Addition 2 $q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha)$ $(2y_{p}\lambda-3x_{p}^{2})$
10 Complete Addition 3 $q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$ $(\lambda^{2}-x_{p}-x_{q}-x_{r})$
11 Complete Addition 4 $q_{\text{add}}\cdot x_{p}x_{q}(x_{q}-x_{p})$ $(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
12 Complete Addition 5 $q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$ $(\lambda^{2}-x_{p}-x_{q}-x_{r})$
13 Complete Addition 6 $q_{\text{add}}\cdot x_{p}x_{q}(y_{q}+y_{p})$ $(\lambda(x_{p}-x_{r})-y_{p}-y_{r})$
@@ 17 Complete Addition 10 $q_{\text{add}}\cdot(1-x_{q}\cdot\gamma)(y_{r}-y_{p})$
18 Complete Addition 11 $q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$ $(y_{q}+y_{p})\delta)x_{r}$
19 Complete Addition 12 $q_{\text{add}}\cdot(1-(x_{q}-x_{p})\alpha$ $(y_{q}+y_{p})\delta)y_{r}$
HyperPlonk (HP) Polynomials [6]
20 Vanilla ZeroCheck $(q_{L}w_{1}+q_{R}w_{2}-q_{O}w_{3}$ $q_{M}w_{1}w_{2}+q_{C})f_{r}$
21 Vanilla PermCheck ($\alpha$ is a scalar) $(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}$ $N_{1}N_{2}N_{3}))f_{r}$
22 Jellyfish ZeroCheck $(q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{3}+q_{4}w_{4}$ $q_{M_{1}}w_{1}w_{2}+q_{M_{2}}w_{3}w_{4} +$$q_{H_{1}}w_{1}^{5}+q_{H_{2}}w_{2}^{5}+q_{H_{3}}w_{3}^{5} +$$q_{H_{4}}w_{4}^{5}-q_{O}w_{5} +$$q_{ecc}w_{1}w_{2}w_{3}w_{4}+q_{C})f_{r}$
23 Jellyfish PermCheck ($\alpha$ is a scalar) $(\pi-p_{1}p_{2}+\alpha(\phi D_{1}D_{2}D_{3}\cdot$ $D_{4}D_{5}-N_{1}N_{2}N_{3}N_{4}N_{5}))f_{r}$
24 OpenCheck $y_{1}f_{r_{1}}+y_{2}f_{r_{2}}+y_{3}f_{r_{3}}$ $y_{4}f_{r_{4}}+y_{5}f_{r_{5}}+y_{6}f_{r_{6}}$
Figure 6: Speedups over 4-threaded CPU implementations of SumChecks on the polynomials in Table I. We pick a different optimal design across each bandwidth technology. Utilization across the design points is also shown.
@@ VI-A1 Speedup over various gate types
Table I enumerates the polynomials we use in our evaluation. We first benchmark the performance of SumChecks on each polynomial by extending HyperPlonk’s existing code base [6]. We then run each polynomial on our programmable SumCheck unit. In this evaluation, we seek to optimize both performance and utilization across all polynomials. We can always allocate EEs and PLs to handle the worst-case polynomial degree, however, this would lead to underutilization on low-degree polynomials. In our design space search, we specify an area and bandwidth constraint and then use the following objective function:
$\min_{\text{design}}\;(1-\lambda)\cdot f_{\text{slowdown}}(s_{d,i})+\lambda\cdot(1-f_{\text{util}}(u_{d,i}))$
Here, $s_{d,i}$ refers to the slowdown of the $i^{th}$ polynomial for a given design $d$, relative to the fastest runtime for that polynomial in the area-constrained design space. Similarly, $u_{d,i}$ reflects the utilization for design $d$ on polynomial $i$. $f_{\text{slowdown}}$ and $f_{\text{util}}$ are aggregation functions; we employ geomean for slowdown and arithmetic mean for utilization. $\lambda$ controls the trade-off between the minimizing slowdown and maximizing utilization. For this analysis, we focus on utilization, using $\lambda=0.8$. We compare the runtime against the CPU running each polynomial’s SumCheck on 4 threads, for which we estimate the total core area to be 37 mm2 in 7nm [44, 1, 58] and use that as our area constraint (after scaling from 22nm). Across various bandwidths, we find the optimal design point and report speedups and utilizations in Figure 6. For most designs, 2 EEs and 5 PLs is optimal.
We observe $\sim$ 50% utilization across benchmarks at each bandwidth, due to three main factors: (1) MLE Update units are inactive during the first round of SumCheck, which accounts for about half of total runtime; (2) lower-degree polynomials require fewer product lanes; and (3) repeating MLEs (e.g., $q_{add}$) reduce demand on the update modular multipliers since they are reused across terms. Consequently, when optimizing for utilization, we observe a tendency to limit the number of EEs. Further parallelism and potentially higher utilization can be achieved by mapping multiple terms to EEs when possible, but it would introduce additional complexity in interconnect and increase bandwidth demand. These trade-offs can be explored further in future work. In contrast, product lane modular multipliers are usually fully utilized, provided the polynomial degree and memory bandwidth are high enough. Despite moderate utilization, we get significant speedups over CPU baselines, approaching $1000\times$ at 1 TB/s of bandwidth.
Our SumCheck unit was benchmarked on a “training set” of mostly lower-degree polynomials, which comprise most ZKP protocols today. However, we found that higher-degree (e.g., Jellyfish polynomials) achieved comparable or higher utilization ($\approx 47\%$) than many lower-degree polynomials. This is because additional constituent polynomials and extension products place greater concurrent demand on EEs and product lanes, indicating that the utilization improves with greater polynomial complexity for a given hardware configuration.
VI-A2 High-degree sweep
@@
We analyze the performance scaling of SumChecks as we increase polynomial degree. We sweep polynomial degrees ($d=2\ldots 30$) for $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+q_{c}$, as in HyperPlonk, and measure speedup over CPU. We use the same objective as before and pick a high-performance design under the same area constraints (Figure 7). Our results show that low-degree polynomials require HBM-scale bandwidths to achieve $1000\times$ speedups, while higher degree polynomials can reach similar speedups with DDR5-level bandwidths (e.g., 256 GB/s). Higher degree polynomials in this scenario exhibit more computations on the same amount of data, which is why the disparity in speedups across bandwidths is less compared to low-degree polynomials.
Figure 8: Scheduler-induced SumCheck behavior at fixed BW and PL settings.
At fixed bandwidth, PE count, and EE-per-PE settings, increasing the number of PLs per PE improves performance roughly proportionally. Similar gains are observed when increasing EE count at fixed PLs. Interestingly, as shown in Figure 8, runtime increases in discrete jumps due to the graph decomposition the scheduler uses to map polynomials to EEs (see Figure 2). For example, under 6 EEs, degree-1166 polynomials have 1 node in their schedule graph, while degree-771111 polynomials require 2. Each added node incurs a sharp runtime jump, as it introduces additional products across all MLEs and terms. Within each node cluster, runtime grows more gradually with degree due to early exit optimizations. These trends may help optimize the SumCheck unit when the degree distribution of polynomials is known, by minimizing high-degree outliers that cause runtime spikes.
@@ VI-A3 Comparison with Prior ASICs
Figure 9 details our speedups relative to a prior ASIC, zkSpeed. zkSpeed implements the SumChecks by using a custom unified core that maximizes immediate reuse. We also compare with zkSpeed+, which gives zkSpeed the benefits of pipelining MLE updates directly into the extensions and product computations (as we do in zkPHIRE). We assume the same modular multiplier area and 2 TB/s as zkSpeed. We use the same objective function from Section VI-A1 on the same training set of polynomials in Table I with a roughly iso-zkSpeed area constraint. We pick a 35.24 mm2 design compared to zkSpeed’s 30.8 mm2 SumCheck and MLE Update area. Our design has 4.9 mm2 of SRAM, while zkSpeed used a much larger global SRAM for storing MLEs, hence we believe this comparison is fair.
@@
We find that, for ZeroCheck, because the Jellyfish polynomial is more complex, the $2\times$ is not sufficient to bring speedups over the Vanilla counterpart. However, PermCheck’s increase in complexity is outweighed by a $2\times$ workload reduction. OpenCheck, which is the same between Jellyfish and Vanilla, exhibits $1.12\times,3.54\times$ and $6.22\times$ speedup for each gate reduction over zkSpeed+. Combining all three SumChecks, a Jellyfish reduction by $4\times$ is sufficient to outperform Vanilla polynomials on both zkSpeed+ and zkPHIRE. These gains are observed in spite of the inherent overheads of programmability in our accelerator.
@@
VI-A4 Comparison with GPU
We evaluate zkPHIRE against the GPU library ICICLE [21] by running SumChecks from the Spartan and HyperPlonk protocols on an NVIDIA A100 GPU. We use the same SumCheck design from Section VI-A3, assuming 1 TB/s of bandwidth to roughly match the A100. Spartan is chosen due to its adoption in several recent protocols [3, 30, 29], while HyperPlonk evaluates performance on high-degree gates. For context, we also report CPU runtimes for these SumChecks in Table II. Notably, ICICLE currently supports composite polynomials with at most eight unique constituent polynomials, preventing evaluation of polynomials 21-24 in Table I. We choose to evaluate the Vanilla Gate portion of polynomial 20 (excluding $f_{r}$). We find that zkPHIRE accelerates these SumChecks by roughly $70\times$ over GPU, and by $600-800\times$ over CPU.
Table II: SumCheck Runtimes on CPU, GPU, and zkPHIRE for $N=24$. HyperPlonk polynomials are from Table I. zkPHIRE runtimes are accompanied with speedups over CPU/GPU
Polynomial Number of Problem Runtime (ms)
Type SumChecks Size CPU GPU zkPHIRE
$(A\cdot B-C)\cdot f_{\tau}$ 1 $2^{N}$ 6770 571 7.6 ($891\times$$75\times$)
$(\text{Sum}_{ABC})\cdot Z$ 1 $2^{N+1}$ 5237 586 8.4 ($623\times$$70\times$)
$A\cdot B\cdot C$ 12 $2^{N}$ 60993 5376 78.9 ($773\times$$68\times$)
$A\cdot B\cdot C$ 6 $2^{N-1}$ 15248 1440 19.7 ($774\times$$73\times$)
$A\cdot B\cdot C$ 4 $2^{N+1}$ 40662 3460 52.6 ($773\times$$66\times$)
[…]
HP Poly. 21 1 $2^{N}$ 21625 22.7 $(953\times)$
HP Poly. 22 1 $2^{N}$ 74226 69.5 $(1068\times)$
HP Poly. 23 1 $2^{N}$ 32774 32.2 $(1018\times)$
HP Poly. 24 1 $2^{N}$ 17591 21.3 $(826\times)$